OpenStudy (anonymous):

A waffle recipe calls for different amounts of mix, depending on the number of servings. For 4 servings, Jordan needs 3 cups of mix and for 12 servings; she needs 9 cups of mix. This shows a direct variation between the number of servings (x) and the amount of mix (y). Using the formula y =kx, solve the formula for k and then give the value of k.

OpenStudy (anonymous):

I updated the last question. the answer is 14mm

OpenStudy (anonymous):

with the formula being 2a/b =h. .. the formula for this one is \[\frac{y}{x}] =k

OpenStudy (anonymous):

\[\frac{y}{x} = k\]

OpenStudy (anonymous):

thanks LOOOOVE (: